OpenStudy (anonymous):

PLEASE HELP<3 i give medals and love haha Solve 4x + 3y = 13 by the substitution method. x + y = 4

OpenStudy (bahrom7893):

igba, your turn to get it right..

OpenStudy (bahrom7893):

x = 4-y from the 2nd equation: 4*(4-y) + 3y = 13

OpenStudy (bahrom7893):

16-4y+3y=13 -y = 13-16 y = 3

OpenStudy (bahrom7893):

x=1

OpenStudy (anonymous):

You guys are life savvvvvers

OpenStudy (anonymous):

y=4-x hene 4x+3(4-x)=13 4x+12-3x=13 x=1

OpenStudy (anonymous):

hence y=3

OpenStudy (anonymous):

@bahrom ...lol

OpenStudy (anonymous):

16-4y+3y=13 -y = 13-16 y = 3 x=1

OpenStudy (anonymous):

PRAISE BAHROM

OpenStudy (anonymous):

You are. Lol

OpenStudy (anonymous):

@bahrom....'on...itz jst dat u saw the question first....

OpenStudy (bahrom7893):

lol

OpenStudy (anonymous):

Shiv is right..

OpenStudy (lgbasallote):

you're actually a legend....of laughs :))) your LoL